A steam turbine is supplied with superheated steam at a pressure of 20 bar and a temperature of 350 °C. The steam exhausts to a condenser at a pressure of 0.1 bar and dryness fraction of 0.85. If the steam flow rate is 2 tonnes per hour and the power output is 420 kW, calculate the thermal efficiency of the energy conversion process.
1
Expert's answer
2018-08-14T10:16:08-0400
h_1=3135.5 kJ/kg
h_2=2225.7 kJ/kg
η=P/(h_2-h_1 ) dm/dt
η=420/(3135.5-2225.7) (2000/3600)=0.256 or 25.6%.